OpenStudy (anonymous):

Which is a solution to the equation y=3x+1 ? A. (3, 10) B. (2, 6) C. (1, 5) D. (0, 3)

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@iGreen

TheSmartOne (thesmartone):

Plug in each answer choice and see if it works.

TheSmartOne (thesmartone):

For example, Let’s plug in Answer Choice D (0,3) that is in the form of (x,y) so: y=3x+1 3=3(0)+1 3=0+1 3 \(\neq\) 1 So D is not correct.